A. A mass of 5kg is on too of a building of height 10m. Calculate its potential energy.
B. The mass then falls at a speed of 8m/s in 4 secs, find its KE and momentum p. What is the force and impulse the mass transfers to the ground?
1
Expert's answer
2019-11-24T16:53:11-0500

"P.E.=mgh=5\times10\times 10=500 joule"

"K.E.=0.5\times m\times v^2"

"K.E=0.5\times 5\times\ 8^2=160joule"

"Force=mg=5\times 10=50tons"

"Impulse= force\times \Delta t=50\times 4=200 Ns""Momentum=mv=5\times 8=40kgm/s"